Let 𝑓: (0,1) β†’ ℝ be a bounded continuous function. Show that 𝑔(π‘₯) = π‘₯(1 βˆ’ π‘₯)𝑓(π‘₯) is uniformly continuous.Β 


1
Expert's answer
2021-12-16T16:04:49-0500

"|g(x)-g(x_0)|=|x(1-x)f(x)-x_0(1-x_0)f(x_0)|="


"=|x(1-x)f(x)-x(1-x)f(x_0)+x(1-x)f(x_0)-x_0(1-x_0)f(x_0)|\le"


"\le |x(1-x)||f(x)-f(x_0)|+|fx_0||x(1-x)-x_0(1-x_0)|"


Since f(x) is bounded we have that "f(x)\le M" where M is some number. Then:

"|x(1-x)||f(x)-f(x_0)|+|fx_0||x(1-x)-x_0(1-x_0)|\le"


"\le 2M|x(1-x)|+M|x(1-x)-x_0(1-x_0)|"


find maximum of "x(1-x)" :

"x=1\/2"

then:

"2M|x(1-x)|+M|x(1-x)-x_0(1-x_0)|<2M\cdot1/4+M\cdot1/4=3M/4"

"\delta=4\varepsilon/3M"

Since "\delta" does not depend on x0, g(x) is uniformly continuous.